majestic aircraft carrier is a manufacturer of single engine high performance turboprop aircraft. MAC is considering purchasing composite wing fixtures for the assembly of its signature aircraft. The cost of the assembly system is $3 million with life expectancy of 10 years, annual operating cost of $200000 with zero salvage value. MAC anticipates a stream of $1000000 in revenue per year for 10 years from this project. Assuming MAC uses a MARR of 10 percent, and e=14% is this investment economically justified ? calculate the ERR.


1
Expert's answer
2021-05-10T14:29:42-0400

Given (external reinvestment interest rate) = 14% and,

MARR = 10%


This project's cash flow diagram will be:





To calculate ERR, we will evaluate the outflows at the start of the project and the inflows at the end of the project.


Value of outflows at the start of the project: "= \$3 million + \$200,000 (P/A, 14\%,10)"

"= \$3,000,000 + \$200,000 (5.229) = \$4,045,800"

At the completion of the project, the amount of inflows is: 

"= \$1,000,000(F/A, 14\% 10)"

"= \$1,000,000(19.385)"

"= \$19.385 million"

Now calculate i(EER), 

"\$4,045,800(1 + i)= \$19.385 million"

"(1 + i) = 4.791"

"(1+ i)= 1.1696"

"16.96\%" is the value of i.

As a result, the ERR is 16.96 percent.

Since the ERR is greater than the MARR, the expenditure is economically justified.